### v1

# chessformer.bib

Abstract
Transformer models have demonstrated impressive capabilities when trained at scale, excelling at difficult cognitive tasks requiring complex reasoning and rational decision-making. In this paper, we explore the application of transformer models to chess, focusing on the critical role of the position encoding within the attention mechanism. We show that in chess, transformers endowed with a sufficiently versatile position encoding can match existing chess-playing models at a fraction of the computational cost. Our architecture significantly outperforms AlphaZero at 8x fewer FLOPS and matches prior grandmaster-level transformer-based agents at 30x fewer FLOPS.
1 Introduction
Chess has long been regarded as a proving ground for advancing artificial intelligence (AI). The importance of strategic planning and rational decision-making in strong chess play make the game an ideal domain in which to test systems aiming to master these human-like cognitive capabilities.
Traditional chess engines employ a specialized tree-search algorithm paired with a handcrafted evaluation function. Most modern chess engines follow this recipe but use an efficiently updated neural network (NNUE) rather than a handcrafted evaluation function.
The AlphaZero engine [AlphaZero] introduced a different recipe based on Monte Carlo Tree Search (MCTS) and deep neural networks. The open-source Leela Chess Zero (Lc0) recreated this approach, building on the same principles with several subsequent enhancements. These engines use large networks that predict not only a position evaluation but also a policy distribution over subsequent moves which is used to guide the search process.
AlphaZero used a convolution-based residual network, which was the leading architecture of the time. However, the state of the art in AI has since shifted from convolution-based residual networks to attention-based transformer models. Transformers resolve the issue of small receptive fields that limit convolution-based architectures and demonstrate "LLM-like scaling", exhibiting impressive capabilities when trained at scale.
We demonstrate that the effectiveness of transformers in chess depends heavily on the choice of position encoding in the attention mechanism. Based on this observation, we adopt the versatile position encoding scheme of Shaw et al. [shaw2018selfattention] and train models endowed with this technique and other enhancements at scale, calling the resulting architecture the ChessFormer. This architecture substantially outperforms prior work in playing strength and puzzle-solving ability at a fraction of the computational cost.
Our contributions are as follows:
  • We present a simple yet performant “ChessFormer” architecture for transformers in chess and apply this architecture across model scales.
  • We ablate the position encoding of our architecture to motivate design choices.
  • We provide a detailed comparison of our models against prior work to demonstrate the superiority of our techniques.
  • We open-source our training code at https://github.com/Ergodice/lczero-training.
This paper is organized as follows. Section 2 describes the vanilla self-attention formulation and compare three position encodings of varying versatility. Section 3 describes our training setup. We compare the results of our models to prior work and run ablation experiments in Section 4. Related work is reviewed in Section 5, and concluding remarks are presented in Section 6.
Figure 1: Elo strength comparison of agents constructed from our CF-6M and CF-240M models against prior art. Our evaluation methodology is described in Section 4.
2 Background
2.1 Self-Attention
Given a sequence of tokens $(\mathbf{x}_{1},\ldots\mathbf{x}_{n})$ where $\mathbf{x}_{i}\in\mathbb{R}^{d}$, self-attention returns a sequence $(\mathbf{z}_{1},\ldots\mathbf{z}_{n})$ where $\mathbf{z}_{i}\in\mathbb{R}^{d}$. Using projection matrices $W^{Q}$, $W^{V}$, and $W^{K}$ in $\mathbb{R}^{d\times d}$, a logit $e_{ij}$ for each pair of tokens $(i,j)$ is computed using scaled dot-product attention:
$e_{ij}=\frac{(\mathbf{x}_{i}W^{Q})(\mathbf{x}_{j}W^{K})^{T}}{\sqrt{d}}$ (1)
The attention weights are obtained via softmax:
$\alpha_{ij}=\frac{\exp(e_{ij})}{\sum_{k=1}^{n}\exp(e_{ik})}$ (2)
Finally, the output is computed as a weighted sum of value projections:
$\mathbf{z}_{i}=\sum_{j=1}^{n}\alpha_{ij}(\mathbf{x}_{j}W^{V})$ (3)
This process is repeated for each head in the self-attention layer, and the outputs are concatenated before being passed through a final linear projection.
2.2 Position Encodings
Because self-attention is permutation-invariant, positional information must be introduced to the model through some kind of position encoding. Many position encoding techniques, like rotary position embeddings [rope], are designed to decay the attention values with the Euclidean distance between tokens.
While Euclidean distance is a useful inductive bias for language and vision applications, we conjecture that it is not sufficient to capture the topology of the chessboard. A square with a bishop, for example, might need to attend to squares on the same diagonal rather than squares that are nearby. To test this, we compare several position encodings. Changes to the vanilla self-attention formulation described in Section 2.1 are marked in blue.
The simplest form of position encoding, the absolute position embedding, adds a possibly learnable bias to each token before the attention layer. Calling these biases $(\mathbf{c}_{1},\ldots\mathbf{c}_{n})$, the absolute position embedding modifies the formula for attention logits as follows:
$e_{ij}=\frac{((\mathbf{x}_{i}{\color[rgb]{0,0,1}+\mathbf{c}_{i}})W^{Q})((\mathbf{x}_{j}{\color[rgb]{0,0,1}+\mathbf{c}_{j}})W^{K})^{T}}{\sqrt{d}}$ (4)
Unlike absolute position embeddings, relative position encodings model the positions of tokens relative to each other. One simple variant introduces relative biases $d_{ij}$ which are added to the attention logits:
$e_{ij}=\frac{(\mathbf{x}_{i}W^{Q})(\mathbf{x}_{j}W^{K})^{T}}{\sqrt{d}}{\color[rgb]{0,0,1}+d_{ij}}$ (5)
For one-dimensional inputs, these bias terms are shared among pairs of tokens $(i,j)$ and $(k,l)$ for which $i-j=k-l$. We adopt the two-dimensional analog of this technique, where two pairs of tokens share a relative position encoding if their horizontal and vertical displacements are the same.
A more general encoding introduced by Shaw et al. [shaw2018selfattention], which we adopt in our final architecture, models the positional relationship between tokens $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ by introducing learnable vectors $a_{ij}^{Q}$, $a_{ij}^{K}$, and $a_{ij}^{V}$ in $\mathbb{R}^{d}$. The calculation for attention logits is altered to:
$e_{ij}=\frac{(\mathbf{x}_{i}W^{Q}{\color[rgb]{0,0,1}+a_{ij}^{Q}})(\mathbf{x}_{j}W^{K}{\color[rgb]{0,0,1}+a_{ij}^{K}})^{T}}{\sqrt{d}}.$ (6)
The output is updated to propagate $a_{ij}^{V}$ in addition to the output of the value projection:
$\mathbf{z}_{i}=\sum_{j=1}^{n}\alpha_{ij}(\mathbf{x}_{j}W^{V}{\color[rgb]{0,0,1}+a_{ij}^{V}}).$ (7)
While this encoding is computationally expensive at large token counts, our models have a context length of 64, making the additional cost negligible.
3 Experimental Setup
We train all our models in the supervised setting on selfplay data generated via the AlphaZero process. The selfplay setup uses a transformer model of roughly 100 million parameters at 600 nodes per move. This model is almost identical in architecture to our main runs, although it uses a slightly weaker position encoding. All models and ablations of the ChessFormer architecture employ a standard encoder-only backbone with context length 64 with one of the position encodings described in Section 2. We describe the architecture in more detail in Appendix A and note other techniques which did not improve performance in Appendix D.
All models were trained with the Nadam optimizer with $\beta_{1}=0.9$, $\beta=0.98$, and $\epsilon=10^{-7}$, and gradient clipping was set to 10. The final model checkpoints we use for evaluation were generated with stochastic weight averaging [swa].
Our largest model, CF-240M, has 15 encoder layers with an embedding depth of 1024, 32 heads per layer, and a feedforward depth of 4096, for a total of 243 million parameters. It was trained for 3.7 million steps with a batch size of 4096 on 8 A100 GPUs using data parallelism. The learning rate was initialized at $1*10^{-3}$ and manually reduced to $3*10^{-4}$ at 3.2 million steps and $1*10^{-4}$ at 3.6 million steps. The dataset consisted of 500 million games generated from mid-2023 to mid-2024, with each game containing roughly 200 positions.
We also train a smaller model, CF-6M, with 8 encoder layers, an embedding depth of 256, 8 heads per layer, and a feedforward depth of 256, for a total of 6 million parameters. The same configuration is used for ablation runs. CF-6M and its ablations were each trained on a single A100 GPU with a batch size of 2048. The learning rate was initialized at $5*10^{-4}$ and reduced to $1.58*10^{-4}$ and $5*10^{-5}$ at 1.6 and 1.8 million steps, respectively. The dataset consisted of 53 million games generated in April 2024.
Our training targets consist of a mix of auxiliary policy and value targets in addition to the main policy and value heads. More details can be found in Appendix B. None of our training runs exhibited overfitting, which is likely due to the size of the dataset relative to the parameter counts of our models.
4 Results
We compare the playing and puzzle-solving abilities of agents constructed from our models to prior work and ablate the position encoding in our architecture to motivate design choices.
We consider two types of agents: those that select the move which is ranked highest in the policy vector predicted by the model, and those that emulate a depth 1 search by evaluating the model for each legal move and selecting the move that maximizes the score. The policy strategy requires a single model evaluation while the score maximization strategy requires an evaluation for each legal move. To estimate the FLOPs used by an agent of the score maximization type, we multiply the model FLOPs by 20, which is roughly equal to the average number of legal moves available in a position.
We construct both types of agents from our models. We denote an agent by its model name followed by “policy” or “value” to indicate whether the agent uses the predicted policy vector or the score maximization approach (e.g., CF-240M-policy). We construct two agents from the AlphaZero model in the same way. Additionally, our analysis includes models from Ruoss et al. [ruoss2024grandmasterlevel] which use the score maximization approach. We use the final checkpoints of their main runs having parameter counts of 9 million, 136 million, and 270 million, which we refer to as GC-9M, GC-136M, and GC-270M, respectively.
4.1 Playing Strength
To estimate the playing strength of our agents, we play 1,000 games with them against the GC-270M agent of Ruoss et al. [ruoss2024grandmasterlevel], tying the Elo rating of GC-270M to the value reported in that paper. We use the same opening book as Ruoss et al. to achieve a similar comparison to theirs. Elo values for the GC and AlphaZero agents are taken from that paper. The results are shown in LABEL:table:main and Figure 1.
4.2 Puzzles
In addition to playing strength, we also evaluate the puzzle-solving ability of our agents using the subset of the Lichess puzzle database curated by Ruoss et al. [ruoss2024grandmasterlevel]. Our evaluation follows the same methodology as theirs, with an agent deemed to solve a puzzle only if it correctly picks each move in the sequence of the solution. The results are shown in LABEL:table:main. In Figure 2, we also graph the accuracy of our two CF-240M agents on this test set against the GC-136M and GC-270M agents of Ruoss et al. [ruoss2024grandmasterlevel].
Figure 2: Puzzle-solving ability on puzzles rated 1000-3000 of our CF-240M-policy and CF-240M-value agents against the GC-136M and GC-270M agents of Ruoss et al. [ruoss2024grandmasterlevel].
4.3 Ablations
We train ablations of the CF-6M model, comparing the three position encodings described in Section 2. As shown in LABEL:table:ablate, our chosen encoding, that of Shaw et al. [shaw2018selfattention], substantially outperforms both the relative bias and absolute position encodings. The reported value loss and accuracy statistics are those of the main “result” value head. At this parameter count, we have observed that doubling model size yields an increase in policy accuracy of around 1.5%.
4.4 Analysis
Our main results can be found in LABEL:table:main. Our agents consistently outperform prior work in both playing strength and puzzle-solving ability at a fraction of the computational cost. Our CF-6M-policy agent outperforms the AlphaZero-policy agent in both puzzle-solving ability and Elo performance at 8x fewer FLOPS. At 30x fewer FLOPS, our CF-240M-policy agent matches the puzzle accuracy and exceeds the Elo performance of the grandmaster-level agent GC-270M of Ruoss et al. [ruoss2024grandmasterlevel].
5 Related Work
Independent work in the same vein as this paper by Ruoss et al. [ruoss2024grandmasterlevel] analyzes the effect of training transformer models at scale for chess, demonstrating that a transformer-based approximation of a strong oracle can be trained without domain-specific alterations to achieve grandmaster-level strength without search. Unlike this paper, they focus on distilling a search-based algorithm into a transformer model rather than optimizing for playing strength.
Focusing instead on maximizing playing strength, Czech et al. [czech2023representation] design a lightweight transformer block for chess and show that the performance of neural chess models can be improved with carefully chosen input representations and value targets. Similar work in Go [wu2020accelerating] adds auxiliary training targets and enhanced input representations to improve convergence speed and final playing strength of the AlphaZero process.
Other work interprets the inner workings of neural chess models through ablations, probing of internal states, and analysis of attention maps. McGrath et al. [mcgrath2021acquisition] analyze the acquisition of chess knowledge by AlphaZero during training, while Jenner et al. [lookahead] examine the role of attention maps and lookahead in an older transformer model trained by the Lc0 project.
6 Conclusion
One key takeaway of this work is that, though the vanilla transformer is an effective generalist architecture, especially when trained at scale, there can be substantial benefit to exploring domain-specific enhancements and inductive biases. This holds true when resources are limited but especially in search-based applications where speed and accuracy are fungible. This idea is exemplified by work such as AlphaFold [alphafold], which predicts the structure of proteins by modeling them with a graph where edges correspond to relationships between residues.
7 Acknowledgements
This paper relied on extensive contributions from the Lc0 community. We are grateful to Twitch user KittenKaboodle for generously providing the hardware on which experiments were run. We are also thankful to David Elliott, Hunter Monroe, and Evan Engler for proofreading and providing insight. Any errors that remain are our own.
[…]
Appendix A Model Architecture
A.1 Input Encoding
The input to our network is a sequence of 64 tokens of length 112 consisting of a concatenation of:
  • 8 one-hot vectors of length 12 describing the piece at that square for the current position and past 7 positions.
  • En passant and castling information.
  • The number of positions since the last capture, pawn move, or castle, divided by 100.
  • Whether each of the current and past 7 positions is a repetition.
To generate token embeddings, we use a linear projection followed by adding and multiplying by learned offset vectors which are separate across tokens and depth. This gives the model absolute positional information and was found in initial experiments to incrementally improve model quality.
A.2 Body
The body of our models consists of a stack of encoder layers with Post-LN normalization and the initialization/gain scheme Deepnet [wang2022deepnet]. We use a fixed context length of 64, i.e., one token for each square on the board read from left to right then bottom to top, and the board is flipped with the side to move. We use Mish [misra2019mish] activations in the feedforward sublayer.
Following [palm], we omit biases in the QKV projections and omit centering and biases in the encoder normalization layers, finding this to increase training throughput without degrading quality. However, we retain biases in the feedforward and post-attention projections since removing them was found to degrade quality in initial experiments.
A.3 Output Heads
Similar to AlphaZero, our models have heads for policy and value predictions. However, we add several auxiliary policy and value targets to increase convergence speed, following research in Go by Wu [wu2020accelerating], [wu2024].
Policy Heads We introduce a new attention policy head based on scaled dot-product attention. Moves are encoded by the starting square of the piece moved and the destination square to which it moved. Given the sequence of tokens outputted by the body, we generate policy embeddings by applying a dense layer of depth equal to model’s embedding size. From this we generate via linear projection a set of query vectors corresponding to the starting square and a set of key vectors corresponding to the destination square, both with depth equal to the depth of the encoder body.
Logits for moves are calculated via standard scaled dot-product attention, described in Equation 1. We did not find improvement from more complicated self-attention variants, likely because the embedding size is large compared to the token count. To improve training stability, illegal moves are always masked.
Value Heads To model value information, we apply a linear projection of depth $d_{\text{value}}$ to the output of the body, where $d_{\text{value}}=32$ by default. We then flatten the result and apply another projection of size 128, which we call the value embedding. Each of our models has three value heads, each of which generates its own value embedding separately from the body output. Each of these value heads uses this embedding to predict one or more training targets.
The “result” value head predicts the result of the game from among “win”, “draw”, and “loss” with cross entropy loss. The “q” and “short-term” value heads each predict three targets: a reward trained with L2 loss, a categorical distribution over rewards trained with cross-entropy loss, and the error between the predicted and true reward. Respectively, these latter two heads use the reward estimate produced during selfplay and an exponential moving average of future rewards with expected depth 6. The gradient of the reward prediction is detached when calculating the error loss to prevent the error head from affecting the reward prediction.
Appendix B Train Losses
As described above, our models train on several auxiliary targets, which were found by the Katago team [wu2020accelerating], [wu2024] to accelerate model convergence and improve final performance. Here we describe the losses for these targets. The final loss function is a weighted sum of these terms.
Policy Targets
  • Vanilla policy head: predicts the policy target produced by self-play.
    $-c_{\text{pol}}\sum_{m\in\text{moves}}\pi(m)\log(\hat{\pi}(m))$
    where $\pi$ is the policy target and $\hat{\pi}$ is the model’s prediction of $\pi$ and $c_{\text{pol}}=1$.
  • Soft policy head: predicts a high-temperature version of the policy target.
    $-c_{\text{softpol}}\sum_{m\in\text{moves}}\pi_{\text{soft}}(m)\log(\hat{\pi}_{\text{soft}}(m)),$
    where $\pi_{\text{soft}}$ is taken from $\pi$ by setting the temperature to 4 and $\hat{\pi}_{\text{soft}}$ is the model’s prediction of $\pi_{\text{soft}}$. We set $c_{\text{softpol}}=8$, though this term is still small compared to the main policy loss because the loss is muted by the high temperature.
Value Targets
  • Game result: predicts the outcome of the game.
    $c_{\text{value-wdl}}\sum_{r\in\{\text{win, loss, draw}\}}v(r)\log(\hat{v}(r))$
    where $v$ is the game result, $\hat{v}$ is the model’s prediction of $v$, and $c_{\text{value-wdl}}=1$.
  • L2 value: predicts a scalar reward produced by selfplay.
    $c_{\text{value-l2}}(q-\hat{q})^{2}$
    where $q$ is the true reward predicted by $\hat{q}$ and $c_{\text{value-l2}}=1$.
  • Categorical value: predicts a categorical distribution over reward estimates.
    $c_{\text{value-cat}}\sum_{x\in\text{score buckets}}z(x)\log(\hat{z}(x))$
    where $z$ is a one-hot vector representing the bucket into which the recorded reward $q$ falls into. We set $c_{\text{value-cat}}=0.1$ for CF-240M and excluded this target for all other training runs.
  • Value error: predicts the error of the L2 value prediction.
    $c_{\text{value-error}}(\hat{e}-(q-\hat{q})^{2})^{2}$
    where $c_{\text{value-error}}$ = 1 and $\hat{e}$ is the model’s prediction of the squared difference between the predicted and true reward.
Appendix C Playstyle
We notice that our largest model, CF-240M, has a very humanlike understanding of the game, which allows it to assess positions in ways that elude top minimax-based engines. One such position arose in a game between Lc0 and Stockfish at the 23rd Computer Chess Championship Rapid tournament (see here), where the former used a slightly older transformer model with 190 million parameters. Lc0 managed to construct a fortress from a lost position, leading the game into a draw. We show a position from this game in Figure 3(a). CF-240M assigns the only drawing move, Kg7, a policy of 92.8%, and evaluates the position as drawn (2.2%W, 76.3%D, 21.5%L). Stockfish assigned the same position a +7 evaluation at 11 billion nodes, and its evaluation stayed above +3 for over 100 moves.
Another position demonstrating this humanlike positional understanding arose in the 21st Computer Chess Championship Rapid tournament (see here), where Lc0 used a medium-sized transformer model of around 100 million parameters. Lc0 played b4, trapping the white rook on b3. CF-240M assigns this move the highest policy at 37% and evaluates the position as slightly winning (40.3%W, 54.1%D, 5.6%L).
If the pawn on b2 is removed and this trapping idea is no longer feasible, the evaluation becomes much more drawish (17.1%W, 77.7%D, 5.2%L), despite black gaining a material advantage. Additionally, the policy on b4 is reduced to 4.76%, and Rc7 becomes the top policy move. Lc0 ended up winning the game, with Stockfish’s evaluation only going negative several moves later despite searching hundreds of millions of nodes.
Our models’ positional understanding seems to extend to long-term planning. As shown in Figure 3(c), CF-240M picks the right move at each of white’s turns in a famous king walk played by GM Nigel Short as white against GM Jan Timman. The line continues 31. Kh2 Rc8 32. Kg3 Rce8 33. Kf4 Bc8 34. Kg5.
(a) Fortress detection: CF-240M finds the only drawing move, Kg7, and evaluates the position as drawn.
(b) Trapped piece detection: when white has a pawn on b2, CF-240M opts to trap the white rook on b3 by pushing b4. However, if the pawn on b2 is not present, it abandons this idea.
(c) Long-term planning: CF-240M finds a famous king walk that was played out by Nigel Short.
Figure 3: Humanlike play: our transformer models exhibit a humanlike understanding of the game, detecting positional ideas that elude top minimax-based engines.
Appendix D Other Techniques Tried
In addition to the techniques reported here, we tried several architectural modifications that did not bear fruit. For example, we found replacing the feed-forward layer with a Gated Linear Unit [glu] to slightly degrade performance when model FLOPS were held constant. Using a sparsely gated [moe] or soft [softmoe] mixture of experts in place of the feed-forward layer also did not improve performance.